# Test-Time Alignment of Discrete Diffusion Models with Sequential Monte Carlo

### v1

@@ Appendix B Sampling Algorithm
Appendix C Gradient analysis of the Denoising model
@@
Appendix D Related Work
SMC. SMC has been proposed for guidance or alignment of continuous diffusion models by Wu et al. (2023); Dou & Song (2024); Cardoso et al. (2023); Phillips et al. (2024); Kim et al. (2025). In the context of discrete diffusion models, Lee et al. (2025) use SMC to sample from the unnormalized tempered distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$, where $\zeta$ is a conditioning variable and $\alpha$ is the guidance scale. They show that guided rate matrix $R_{t}^{\alpha}$ normally used for guidance based sampling is equal to the true tempered rate matrix $R_{t}^{\alpha,\text{true}}$ only for $\alpha=1$. Thus, to sample from $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$ where $\alpha\neq 1$, they propose to decouple the rate matrix used as $R_{t}^{\beta}$ from $\alpha$. SMC is used to sample from the target distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$, while the distributions induced by the rate matrix $R_{t}^{\beta}$ serve as the proposal. Usually, $\beta$ is set close to 1. In contrast to our method, their method necessitates additional learning or fine-tuning to first obtain the rate guided rate matrix $R_{t}^{\alpha=1}$ which is needed for the proposal. Additionally, the proposal used is not necessarily optimal to minimize weight variance. As $\alpha$ and $\beta$ diverge, SMC resampling will be doing most of the heavy-lifting (instead of the guidance) and this may lead to high resampling frequency and low sample diversity.
@@ Appendix D Related Work
Appendix E Experimental Setting and Additional Results
E.1 Baselines
@@
E.2 Mixture of Gaussian datasets
For all experiments on these datasets, we use 2000 particles and 100 discretized time steps. The reward function computation details are provided in the Sec. E.2.1. For resampling, we have used ESS threshold $\text{ESS}_{\text{min}}=1000$, and we perform a full resample using systematic multinomial sampling. The KL weight $\alpha$ is 1, and the temperature schedule is set as $\lambda_{t}=1-t$. For the diffusion model, we use a linear noise schedule $\alpha_{t}=1-t$.
E.2.1 Reward computation
@@
E.2.2 Reward and ESS traces for different proposals
@@
E.3 Binarized MNIST dataset
@@ E.3.2 Using UDLM
Appendix F Limitations and Future Work
While our method shows good results on small datasets, its effectiveness on more complex tasks, such as math reasoning, protein design, remains to be evaluated. Additionally, we have yet to benchmark our approach against a broader set of reinforcement learning and guidance methods, as outlined in Appendix D. The temperature schedule $\lambda_{t}$ used for intermediate target distributions is presently hand-crafted; future work will explore automating this component through adaptive tempering techniques. We also see potential in extending the method to continuous-time discrete diffusion models (Campbell et al., 2022; Sun et al., 2022), which may further enhance its flexibility and applicability.

### v2

@@ Appendix C Sampling Algorithm
Appendix D Gradient analysis of the Denoising model
@@
Appendix E Related Work
SMC. SMC has been proposed for guidance or alignment of diffusion models by Wu et al. (2023); Dou & Song (2024); Cardoso et al. (2023); Phillips et al. (2024); Kim et al. (2025); Singhal et al. (2025); Skreta et al. (2025); He et al. (2025). In the context of discrete diffusion models, Lee et al. (2025) use SMC to sample from the unnormalized tempered distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$, where $\zeta$ is a conditioning variable and $\alpha$ is the guidance scale. They show that guided rate matrix $R_{t}^{\alpha}$ normally used for guidance based sampling is equal to the true tempered rate matrix $R_{t}^{\alpha,\text{true}}$ only for $\alpha=1$. Thus, to sample from $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$ where $\alpha\neq 1$, they propose to decouple the rate matrix used as $R_{t}^{\beta}$ from $\alpha$. SMC is used to sample from the target distribution $p_{0}(x_{0})p_{0}(\zeta|x_{0})^{\alpha}$, while the distributions induced by the rate matrix $R_{t}^{\beta}$ serve as the proposal. Usually, $\beta$ is set close to 1. In contrast to our method, their method necessitates additional learning or fine-tuning to first obtain the rate guided rate matrix $R_{t}^{\alpha=1}$ which is needed for the proposal. Additionally, the proposal used is not necessarily optimal to minimize weight variance. As $\alpha$ and $\beta$ diverge, SMC resampling will be doing most of the heavy-lifting (instead of the guidance) and this may lead to high resampling frequency and low sample diversity.
@@ Appendix E Related Work
Appendix F Experimental Setting and Additional Results
F.1 Baselines
@@
F.2 Mixture of Gaussian datasets
For all experiments on these datasets, we use 2000 particles and 100 discretized time steps. The reward function computation details are provided in the Sec. F.2.1. For resampling, we have used ESS threshold $\text{ESS}_{\text{min}}=1000$, and we perform a full resample using systematic multinomial sampling. The KL weight $\alpha$ is 1, and the temperature schedule is set as $\lambda_{t}=1-t$. For the diffusion model, we use a linear noise schedule $\alpha_{t}=1-t$.
F.2.1 Reward computation
@@
F.2.2 Reward and ESS traces for different proposals
@@
F.3 Binarized MNIST dataset
@@ F.3.2 Using UDLM
Appendix G Limitations and Future Work
While our method shows good results on small datasets, its effectiveness on more complex tasks, such as math reasoning, protein design, remains to be evaluated. Additionally, we have yet to benchmark our approach against a broader set of reinforcement learning and guidance methods, as outlined in Appendix E. The temperature schedule $\lambda_{t}$ used for intermediate target distributions is presently hand-crafted; future work will explore automating this component through adaptive tempering techniques.